# Yesterday’s News: Benchmarking Multi-Dimensional Out-of-Distribution Generalisation of Misinformation Detection Models

### v1

@@ 2.2 Misinformation Datasets
    Publisher: experts label publishers by their propensity for factual reporting. Article labels are generated by proxy through the label of their publisher Horne et al. (2018); Barrón-Cedeño et al. (2019); Cui and Lee (2020)
Fine-grained annotation methods yield high-quality labels, but are prohibitively expensive to procure. As a result, such datasets usually sample from few publishers, or a relatively limited time-span. Using publisher labels as a proxy can introduce noise, as it assumes all publishers’ articles share a label, but is far cheaper. This can enable large collections of articles, with diverse sets of publishers and events. The ‘Claim’ annotation level, is a relatively new method, and could balance the precision of ‘Span’ annotations with the affordability of ‘Publisher’ labels. misinfo-general is an example of ‘Publisher’ labelling.
We recommend Xiao and Mayer (2024) for a more comprehensive review of datasets at the article and claims labelling scales.
2.2.1 Metadata Provenance
For publisher-level annotations and metadata, we scraped Media Bias/Fact Check [^2] https://mediabiasfactcheck.com/ (MBFC). MBFC is a curated database of news publishers, with thorough analyses of publisher origins, bias, and credibility. Despite being run by lay volunteers, MBFC labels correlate well with professional fact-checking sources Kiesel et al. (2019); Broniatowski et al. (2022); Pratelli and Petrocchi (2022). MBFC labels have been used in many earlier works Rashkin et al. (2017); Baly et al. (2018a, 2020a); Burdisso et al. (2024); Casavantes et al. (2024); Szwoch et al. (2024). MBFC labels are dynamic and annotations can change (although this is infrequent). We use the latest available version of labels (Oct. 2024).
2.3 Publisher Reliability Estimation
The estimation of a publisher’s reliability is a relatively well-studied problem. This is usually achieved through a mix of content-based Rashkin et al. (2017); Bianchi et al. (2024) and metadata features Baly et al. (2018a, b, 2019, 2020a, 2020b); Nakov et al. (2021).
Relative to purely content-based classifiers, publisher-level classification can greatly reduce the computational cost needed for classification Burdisso et al. (2024). However, this usually involves incorporating additional historical context, world-knowledge Yang and Menczer (2024) or social context Pratelli et al. (2024).
This mimics how moderators or users might analyse the reliability of a publisher, potentially before ingesting the article contents. However, such approaches might fail in cases of where publishers are unknown, ambiguous or have changed their editorial bias. In those situations, moderation decisions at the article-level is necessary. While misinfo-general is suited for either approach, we focus on testing the generalisation of inductive article-level classifiers.
3 Data
@@
The original authors’ goal was to study the dynamic behaviour of news and news publishers. Existing corpora were deemed inadequate due to having 1) few, relatively homogeneous articles or publishers, 2) too narrow a focus on specific events, 3) bias towards popular publishers, and 4) limited ground truth labelling Horne et al. (2018); Norregaard et al. (2019).
@@ 3.1 Publisher Labelling
From the 2018 iteration onwards, the NELA datasets come with publisher-level labels. However, due to inconsistencies across dataset iterations and the frequency of labelling errors, we chose to relabel the dataset completely. We first mapped all articles to a consistent set of publishers using the URL domain of the scraped articles before removing any publishers known to be news aggregators or social media sites. After updating the article-publisher mapping, we label each publisher using a more recent version of the MBFC database. We provide further detail in Appendix B.1. Ultimately, we identified $488$ distinct publishers, many of which were falsely attributed in NELA’s original set of publishers. The metadata available for each publisher is provided in Appendix B.6.
@@ 3.2 Data Processing
In the remaining texts, we mask various forms of private or identifiable information (PII), both to enhance safety and reduce the number of available classification ‘shortcuts’. We furthermore standardize the copyright masking procedure introduced in Gruppi et al. (2020). This introduces $4$ new special tokens: <copyright> replacing NELA’s repeated @ tokens, <twitter>, <url> and <selfref> for any self-references.
@@ 3.3 Topic Clustering
One of our aims is to test model generalisation across different events and topics. To discover these, we used a modified variant of BERTopic Grootendorst (2022) with a gte-large [^4] https://huggingface.co/thenlper/gte-large Li et al. (2023) backbone. This produced thousands of event clusters for every dataset iteration, each with a TF-IDF representation vector. We aggregate these events into overarching topics by applying spectral clustering to the adjacency matrix induced by the inter-event cosine similarity of the TF-IDF matrix. We arbitrarily limit the number of topics to 10, each with varying numbers of events in them. This process is further described in Appendix B.4.
This extends the work of Litterer et al. (2023) on identifying ‘news storms’ in the NELA corpora to a larger time-span, and a larger set of publishers.
4 Generalisation Taxonomy
Generalisation Axis In Distribution Out-of-Distribution
Time Time
CNN AP Vox
2018 2018 2018
CNN AP Vox
2017 2020 2019
@@ 4 Generalisation Taxonomy
[…]
Publisher Publisher
CNN MSNBC OANN
Reuters AP True Activist
[…]
AP Reuters Fox News
Centre Centre Right
Vox Daily Beast True Activist
Left Left Left
[…]
Vox NYT OANN
Reliable Reliable Questionable
MSNBC 911Truth.org Age of Autism
Reliable Conspiracy Pseudosci.
Table 1: A schematic overview of the generalisation taxonomy. The left columns provide relevant generalisation category and axis, whereas the right columns provide examples of in domain and out-of-distribution article sets.
Prior work has found misinformation models to be brittle to various forms of distribution shift. We partially attribute this to a lack of diversity in article content and publishers in the training data. By relaxing the assumption of article-level labelling, we can construct a dataset that is both large and diverse, enabling generalisation-focused training and evaluation across various axes. Each such axis tests performance with respect to a specific shift in the training and inference time input distributions of misinformation detectors.
This paper considers 6 specific generalisation axes. Time-based generalisation measures the extent to which changes in publisher style affect a model’s predictions. Evolution of article content will also impact performance. We focus on two specific forms of such change: 1) due to spontaneous events, which we define as news-worthy happenings with a definite time-span, or 2) due to changing topics, which we define as large, overarching collections of events that remain relatively static over a period measured in years or decades. We expect language to be noticeably different across different collections of events or topics.
The distribution of publishers is also expected to change. All publishers exhibit some editorial bias or style which can be memorized by classification models. One, usually implicit, expectation of misinformation detectors is generalisation to different political biases or misinformation types. Predictions ought to be based on a publisher’s intent, not their norms and values.
This taxonomy is by no means exhaustive, and we discuss some additional categories of domain shift in Sec. 8.
@@ 4.1 Data Splits
To operationalise these generalisation axes, we build 6 (+1 baseline) train/test splits of the dataset using the publisher-level metadata available to us. In each scenario, a validation split is sampled i.i.d. from the training set. Throughout, we approximate the same $70$/$10$/$20$% article proportions per training/validation/test split, respectively. We repeat each experiment for each year, independently, 6 times in total. The only exception is the ‘Event’ axis, for which we combine all years into a single dataset.
These splits are carefully constructed to test each form of generalisation in isolation. From the model’s perspective, each experiment seems identical, with similar article counts and class proportions. Briefly, we use the following splits:
@@
    Uniform: standard stratified random splitting of articles into disjoint sets
@@
Time: the training set comprises a single dataset year, while the evaluation set consists of articles from seen publishers in all other years
@@
Event: we reserve all articles which include keywords related to the COVID-19 pandemic for testing, and train on the rest
@@
Topic: we reserve the smallest topic clusters for the test set, and train on articles from the rest
@@
Publisher: we reserve the least frequent publishers for the test set
@@
Political Bias: we reserve all articles from either all leftist or rightist publishers for testing
@@
Misinformation Type: we reserve all articles from either all ‘Questionable Source’ or ‘Conspiracy-Pseudoscience’ publishers for testing
@@
We include an expanded description of each split’s construction in Appendix A.
@@
Conspiracy-Pseudoscience 0.43 0.42 -0.01 0.87 0.82 -0.05 0.53 0.53 0.01
Questionable Source 0.43 0.23 -0.20 0.94 0.62 -0.33 0.41 0.25 -0.16
Table 2: Article-level classification performance for various generalisation axes. Specifically, compares performance on the uniformly split in-domain data (ID) to that of the out-of-distribution evaluation set (OoD). The top row uses uniform splitting for both (OoD is actually ID), serving as a baseline value. We have opted to exclude F1 values for the ’Time’ based splitting, as these compare across datasets with strongly varying class proportions.
5 Experiments and Results
To showcase the utility of misinfo-general for model training and evaluation, we trained a simple yet powerful baseline model. Specifically, we fine-tune an instance of DeBERTa-v3555https://huggingface.co/microsoft/deberta-v3-base He et al. (2023) where we reset the pooler and classification weights but freeze the model’s remaining weights. To enable using dataset-specific tokens, we allow the token embedding layer to train with a very low learning rate. The model’s pre-training data included a closed-source news dataset (CC-News), dated between September 2016 and February 2019 Liu et al. (2019).
We fine-tune the models on different splits outlined in Sec. 4. The validation set is always sampled uniformly at random from the training data. We keep the hyperparameters and compute budget constant (which were tuned on the validation sets of the ‘Uniform’ splits). We outline hyperparameters in Appendix D. Training occurs at the article level. We binarise the article publisher’s MBFC label for training labels, which we optimise for using the cross-entropy loss. All ‘Questionable Source’, ‘Conspiracy-Pseudoscience’, and ‘Satire’ publishers were deemed unreliable-, and all others reliable+. We leave testing other publisher-label mappings to future work.
To assess model performance at the article level, we employ the F1-score computed independently for each class along with the Matthews Correlation Coefficient (MCC). The F1-score’s interpretation is largely dependent on the class proportion, making it less suited to comparison across experiments. MCC is more robust to this.
5.1 OoD Generalisation
Table 2 displays the article level classification results for the various generalisation splits outlined in Sec. 4. The larger the deviation between the in distribution (ID) articles in the validation set and the out-of-distribution (OoD) articles in the test set, the worse we consider the model’s generalisation performance.
Firstly, we note that classification performance falls short of desired. While the F1-score for the reliable+class tends to be high (in the range of $0.85-0.95$ at a $\approx$60class proportion), classifying unreliable-articles is considerably more difficult—a trend that holds consistently across generalisation forms. This is largely due to low recall scores for the unreliable-class.
Both ‘Publisher’ and ‘Topic’ splitting show moderate decreases in MCC scores, carried primarily by a decrease in the F1-scores for the unreliable-class. Generalisation to completely unseen publishers or topics, cases where one would expect distinctly different linguistic style or vocabulary, is more challenging. The magnitude of this performance degradation, however, is smaller than we initially expected. We attribute this to two effects:
@@ 5.1 OoD Generalisation
    More mainstream, prolific publishers are obscuring performance on publishers with fewer articles (see App. C). We correct for this effect by including a publisher-level analysis in Sec. 6.1
  • 2.
    The training data is heterogeneous enough for the models to learn generalisation across publishers. We test for this in Sec. 6.2
  • Since it is conceivable that different publishers prefer particular topics, we compute a correlation degree between the produced test sets. While we find a small but consistent overlap between the two, we do not believe this alone accounts for the similarity in performance (see App. A.2).
    The final two generalisation axes exclude a particular misinformation type or political bias from the training set. For the former, we can see little to no effect when removing the ‘Conspiracy-Pseudoscience’ class of articles, but a drastic one if removing ‘Questionable Source’ articles. We posit this is due to the ‘Questionable Source’ being the class of articles written with the explicit purpose of mimicking reliable+publishers, whereas ‘Conspiracy-Pseudoscience’ tends to discuss completely separate topics. In other words, the conspiracy or pseudo-scientific articles tend to be easier to identify as unreliable-.
    For the ‘Political Bias’ generalisation, we see an inability to generalise to opposing political biases. Training on centre and right biased articles sees a $0.19$ drop in MCC, whereas training on centre and left yields a drastic $0.37$ drop. While this is a form of publisher splitting, in both cases the magnitude of the degradation is substantially larger. Especially for transfer to right-biased articles, there exists a drop for both reliable+and unreliable-classification, indicating that it is more challenging for the model to determine article reliability.
    5.2 Generalisation across time
@@
    colspec=cccccccc, columns = colsep=4pt, cell33 = bg = lightgray, cell44 = bg = lightgray, cell55 = bg = lightgray, cell66 = bg = lightgray, cell77 = bg = lightgray, cell88 = bg = lightgray
    Table 3: Article level MCC scores of models trained with uniform splitting on different years of the dataset. Models are evaluated only on publishers also seen during fine-tuning.
@@
    When applying the models to unseen years, we find the models to be surprisingly robust, as shown in Table 5.2. At the article level, despite consistent degradation in performance, proximal years tend to achieve similar scores. Only on very distant years does performance degrade dramatically.
    We speculate that these differences are due to differences in the various dataset iterations, while publisher style or idiosyncrasies being relatively static. For example, all models not trained on the 2017 iteration perform poorly on the 2017 iteration (between $0.26$ and $0.34$ MCC), whereas the 2020–2022 editions perform reasonably well on each other’s years. Indeed, visually, Table 5.2 correlates strongly with Appendix C Table B.6, measuring the amount of overlap in publishers across dataset years.

### v2

@@ 4 The misinfo-general Dataset
    masking self-references, along with other PII (Section 4.3)
In Section 8.1 we show that these pre-processing steps have made publisher identification from articles alone difficult.
In this section, we describe how we gather the dataset content and labels, and generate any additional metadata. In later sections, using article-level metadata, we specifically test for model overfit to publisher style (Section 5 & Section 6), we show including a diverse set of publishers is beneficial to generalization performance (Section 7.2), and we try to find publishers with high degrees of mislabelling by assessing the necessity of model memorization (Section 8.2).
4.1 Article Provenance
@@
The original authors’ goal was to study the dynamic behavior of news and news publishers. They deemed existing corpora inadequate for their goals, because of (1) a small, relatively homogeneous collection of articles or publishers, (2) too narrow a focus on specific events, (3) bias towards popular publishers, and (4) limited ground truth labelling Horne et al. (2018); Norregaard, Horne, and Adali (2019).
@@ 4.2 Publisher Labelling
From the 2018 iteration onwards, the NELA datasets come with publisher-level labels. However, due to inconsistencies across dataset iterations and the frequency of labelling errors, we chose to relabel the dataset completely.
Similar to the initial NELA corpora labels, we scraped Media Bias/Fact Check [^3] https://mediabiasfactcheck.com/ (MBFC). MBFC is a curated database of news publishers, with thorough analyses of publisher origins, bias, and credibility. Despite being run by lay volunteers, MBFC labels correlate well with professional fact-checking sources Kiesel et al. (2019); Broniatowski et al. (2022); Pratelli and Petrocchi (2022). MBFC labels have been used in many earlier works Rashkin et al. (2017); Baly et al. (2018a, 2020a); Burdisso et al. (2024); Casavantes et al. (2024); Szwoch et al. (2024). MBFC labels are dynamic and annotations can change (although this is infrequent). We use those available as of Oct. 2024, well after the publication dates of articles in the corpus.
Using the URL domain of the scraped articles, We first mapped all articles to a consistent set of publishers before removing any publishers known to be news aggregators or social media sites. This gives an article-publisher mapping that is consistent across dataset iterations, and removes cases of where articles were republished on different sites. Each publisher was linked to a publisher in the scraped MBFC database. We provide further detail in Appendix B.1. Ultimately, we identified $488$ distinct publishers, many of which were falsely attributed in NELA’s original set of publishers. The metadata available for each publisher is provided in Appendix B.6.
@@ 4.3 Data Processing
In the remaining texts, we mask various forms of private or identifiable information (PII), both to enhance safety and reduce the number of available classification ‘shortcuts’. We furthermore standardize the copyright masking procedure introduced in Gruppi, Horne, and Adalı (2020). This introduces $4$ new special tokens: <copyright> replacing NELA’s repeated @ tokens, <twitter>, <url> and <selfref> for any self-references.
@@ 4.4 Topic Clustering
One of our aims is to test model generalization across different events and topics. To discover these, we used a modified variant of BERTopic Grootendorst (2022) with a gte-large [^4] https://huggingface.co/thenlper/gte-large Li et al. (2023) backbone. This produced thousands of event clusters for every dataset iteration, each with a TF-IDF representation vector. We aggregate these events into overarching topics by applying spectral clustering to the adjacency matrix induced by the inter-event cosine similarity of the TF-IDF matrix. We arbitrarily limit the number of topics to 10, each with varying numbers of events in them. This process is further described in Appendix B.4.
This extends the work of Litterer, Jurgens, and Card (2023) on identifying ‘news storms’ in the NELA corpora to a larger time-span, and a larger set of publishers.
5 Generalization Taxonomy
In this section, we describe various dimensions along which we believe covariate shifts likely to occur, and which are feasible to simulate using misinfo-general. We consider a total of $6$ specific generalization axes.
Time based generalization measures the extent to which changes in publisher style affect a model’s predictions. The publishers considered in each split should be held constant to avoid confounding with different publishers.
Evolution of article content will also impact performance. We focus on two specific forms of such change: (1) due to spontaneous events, which we define as news-worthy happenings with a definite and narrow time-span, or (2) due to evolving topics, which we define as large, overarching collections of events that remain relatively static over a long period. Across these events and topics, we expect markedly different language.
The distribution of publishers is also expected to change between training and inference time. All publishers exhibit some form of editorial bias or style, which can be memorized by classification models. While models should use style to inform moderation decisions, they should also not overfit to stylistic idiosyncrasies. One related, usually implicit, expectation of misinformation detectors is a robustness to different political biases or misinformation types. Predictions ought to be based on a publisher’s intent, not their norms and values. By excluding these from training, we can test a model’s ability to generalize to different classes of publishers.
Table 1: A schematic overview of the generalization taxonomy. The left columns provide relevant generalization category and axis, whereas the right columns provide examples of in domain and out-of-distribution article sets.
{tblr}
colspec = @X[1,c]X[1,c]X[3,c]@X[3,c]@, hline1,2,8 = 0.5pt,solid, hline3,5 = 0.5pt,solid,gray, hline4,6,7 = 0.5pt,solid,gray, cell11 = r=1, c=2c, cell31 = r=2, c=1c, cell51 = r=3, c=1c, row5 = 2.0em, Generalisation Axis & In Distribution Out-of-Distribution
Time Time CNN AP Vox 2018 2018 2018 CNN AP Vox 2017 2020 2019
@@ 5 Generalization Taxonomy
Topic Crime, Sports Elections
Publisher Publisher CNN MSNBC OANN Reuters AP True Activist
Political Bias AP Reuters Fox News Centre Centre Right Vox Daily Beast True Activist Left Left Left
Misinfo Type Vox NYT OANN Reliable Reliable Questionable MSNBC 911Truth Age of Autism Reliable Conspiracy Pseudosci.
@@ 5.1 Data Splits
To operationalize these generalization axes, we build 6 (+1 baseline) train/test splits of the dataset using the publisher-level metadata available to us. Each split is meant to simulate one of the above described covariate shift scenarios, while ensuring minimal cross-scenario confounding.
Throughout, we approximate the same $70$/$10$/$20$% article proportions per training/validation/test split, respectively. The validation split, used for early stopping, is sampled i.i.d. from the training set. For all scenarios, we repeat each split independently for each dataset year, for a total of 6 times. The only exception is the ‘Event’ axis, for which we combine all years into a single dataset.
Briefly, we construct splits (schematically displayed in Table 2) for the scenarios as follows:
@@
    Uniform: standard stratified random splitting of articles into disjoint article sets. No article meta-data is used
  2. 1.
@@
    Time: the training set consists of a single dataset year, while the test set contains articles from publishers seen during training in all other dataset years. This tests within publisher variation
  3. 2.
@@
    Event: the dataset has been annotated for several thousands of events, but we focus on a singular one: the COVID-19 pandemic. We reserve all articles containing any related keywords for testing, and we train on all non-COVID articles
  4. 3.
@@
    Topic: we reserve the $k$ smallest topic clusters for the test set, such that these contain roughly ${}20\%$ of all articles, and we train on the remaining articles
  5. 4.
@@
    Publisher: similarly, we reserve the $k$ least frequent publishers for the test set, such that these contain roughly ${}20\%$ of all articles, and we train on the remaining articles
  6. 5.
@@
    Political Bias: we reserve all articles from either all ‘Left’- or ‘Right’-biased publishers for testing, and train on articles from the opposite political bias, along with any ‘Center’-biased publishers
  7. 6.
@@
    Misinformation Type: similarly, we reserve all articles from either all ‘Questionable Source’ or ‘Conspiracy-Pseudoscience’ publishers for testing, and train on articles from the other misinformation class. We use an i.i.d. split of reliable articles to ensure a similar class distribution in all splits
@@
We include a substantially expanded description of each split’s construction in Appendix C.
It is important to note that from the model’s perspective, each scenario seems identical. The same labels are present in each split, with roughly the same article counts in the same class proportions. Without additional context, one should expect similar performance across these splits.
6 Experiments and Results
To showcase the utility of misinfo-general for model training and evaluation, we use a simple yet powerful baseline model. Specifically, we fine-tune an instance of DeBERTa-v3555https://huggingface.co/microsoft/deberta-v3-base He, Gao, and Chen (2023) where we reset the pooler and classification weights but freeze the model’s remaining weights. To enable using dataset-specific tokens, we allow the token embedding layer to train with a very low learning rate. The model’s pre-training data included a closed-source news dataset (CC-News), dated between September 2016 and February 2019 Liu et al. (2019), and thus should be easily adapted to misinfo-general. Similar architectures have shown surprisingly adequate performance on other benchmark datasets, including various NELA versions Pelrine, Danovitch, and Rabbany (2021); Zhou et al. (2021); Raza and Ding (2022).
We fine-tune the models on the different splits outlined in Section 5. We keep the hyperparameters and compute budget constant (which were tuned on the validation sets of the ‘Uniform’ splits), which we outline in Appendix D. Training occurs at the article level, using publisher-level labels. We binarize the article publisher’s MBFC label for training labels: all ‘Questionable Source’, ‘Conspiracy-Pseudoscience’, and ‘Satire’ publishers were deemed unreliable- , and all others reliable+ . Other publisher-label mappings have been used in other works, and is deserving of future research for this dataset.
To assess model performance at the article level, we employ the F1-score computed independently for each class along, with the Matthews Correlation Coefficient (MCC). The F1-score’s interpretation is largely dependent on the class proportion Flach and Kull (2015), making it less suited to comparison across experiments, whereas MCC is more robust to this Chicco and Jurman (2020, 2022). MCC is 0 for random performance, and 1 only for perfect classification.
6.1 OoD Generalization
Table 2 displays the article level classification results for the various generalization splits outlined in Section 5. The larger the deviation between the in distribution (ID) articles in the validation set and the out-of-distribution (OoD) articles in the test set, the worse we consider the model’s generalization performance.
Firstly, we note that classification performance falls short of desired. While the F1-score for the reliable+ class tends to be high (in the range of $\mathbf{0.85-0.95}$ at a $\sim 60$% class proportion), classifying unreliable- articles is considerably more difficult—a trend that holds consistently across generalization forms. This is largely due to low recall scores for the unreliable- class. This is especially surprising given the high accuracy scores reported for similar models on other misinformation datasets.
Both ‘Publisher’ and ‘Topic’ splitting show moderate decreases in MCC scores, carried primarily by a decrease in the F1-scores for the unreliable- class. Generalization to completely unseen publishers or topics, cases where one would expect distinctly different linguistic style or vocabulary, is more challenging. The magnitude of this performance degradation, however, is smaller than we initially expected. We attribute this to two effects:
Table 2: Article-level classification performance comparing performance on the ID and OoD evaluation sets. The top row uses uniform splitting for both (OoD = ID), serving as a baseline value. ‘Time’ based splitting has strongly varying class proportions, making F1 values inappropriate.
@@ 6.1 OoD Generalization
Consp.-PSci. 0.43 0.42 -0.01 0.87 0.82 -0.05 0.53 0.53 0.01
Questionable 0.43 0.23 -0.20 0.94 0.62 -0.33 0.41 0.25 -0.16
@@
    More mainstream, prolific publishers are obscuring performance on publishers with fewer articles (see Appendix B.8). We correct for this effect by including a publisher-level analysis in Section 7.1
  2. 2.
    The training data is heterogeneous enough for the models to learn generalization across publishers. We test for this in Section 7.2
Since it is conceivable that different publishers prefer particular topics, we compute a correlation between the produced test sets. While we find a small but consistent overlap between the ‘Publisher’ and ‘Topic’ test sets, we do not believe this alone accounts for the similarity in performance (see Appendix C.2).
The final two generalization axes exclude a particular misinformation type or political bias from the training set. For the former, we can see little to no effect when removing the ‘Conspiracy-Pseudoscience’ class of articles, but a drastic one if removing ‘Questionable Source’ articles. We posit this is due to the ‘Questionable Source’ being the class of articles written with the explicit purpose of mimicking reliable+ publishers, whereas ‘Conspiracy-Pseudoscience’ tends to discuss completely separate topics. In other words, the conspiracy or pseudo-scientific articles tend to be easier to identify as unreliable- .
For the ‘Political Bias’ generalization axis, we see an inability to generalize to opposing political biases. Training on center and right biased articles sees a $\mathbf{0.19}$ drop in MCC, whereas training on center and left yields a drastic $\mathbf{0.37}$ drop. While this is a form of publisher splitting, in both cases the magnitude of the degradation is substantially larger. Especially for transfer to right-biased articles, there exists a drop for both reliable+ and unreliable- classification, indicating that it is more challenging for the model to determine article reliability.
Table 3: Article level MCC scores of models trained with uniform splitting on different years of the dataset.
@@
2017 2018 2019 2020 2021 2022
[…]
2018 0.29 0.42 0.43 0.39 0.41 0.37
2019 0.26 0.38 0.44 0.40 0.41 0.40
2020 0.34 0.39 0.47 0.47 0.47 0.45
2021 0.31 0.37 0.46 0.46 0.47 0.45
2022 0.33 0.38 0.46 0.45 0.46 0.46
6.2 Generalization across time
When applying the models to unseen years, we find the models to be surprisingly robust, as shown in Table 3. At the article level, despite consistent degradation in performance, proximal years tend to achieve similar scores. Only in very distant years does performance degrade dramatically.
We speculate that these differences are due to differences in the various dataset iterations, while publisher style or idiosyncrasies being relatively static. For example, all models not trained on the 2017 iteration perform poorly on the 2017 iteration (between $\mathbf{0.26}$ and $\mathbf{0.34}$ MCC), whereas the 2020–2022 editions perform reasonably well on each other’s years. Indeed, visually, Table 3 correlates strongly with Appendix B.8 Table B.8, showing the amount of overlap in publishers across dataset years.